No significant change was found inside the vacuum. We still see clipping at the Faraday, and also, we saw clipping by BS coil holder. Using PZT1, we could make it better, but this might be causing PRC problem -- BS is inside the PRC, too.
Yuta just told Jamie and I that when he and Koji were looking at things yesterday, they saw that the beam spot was roughly at the center of the PRM, but was clipping on the lower OSEM holder plate on the BS. This indicates that the beam spot on the BS is much too low. The easiest way I can see this happening is poor pitch pointing with the tip tilts, which we unfortunately don't have active control over.
Recall elog 3425, where I mentioned some pretty bad pitch pointing after a TT was moved from the cleanroom, to the chamber, back to the cleanroom. I think that we may need to check the pitch pointing at the chamber before installing tip tilts in the future.